OpenStudy (anonymous):

Simultaneous equation y= x^2 y + 3x = 4 Anyone can help with the steps to do it?

OpenStudy (llort):

since y=x^2 you can substitute into the second equation x^2+3x-4=0 what adds to +3 and multiplies to -4?

OpenStudy (anonymous):

Start by making both equations equal to y. y + 3x = 4 <=> y = 4 - 3x. Then set the equations equal to each other, and solve for x. x^2 = 4 - 3x x^2 + 3x - 4 = 0 (x - 1)(x + 4) = 0 The equations intersect when x =1, and x = -4, so you put those values into your original equation. y = (1)^2; y = (-4)^2. y = 1; y = 16. The functions intersect at (1,1) and (-4, 16).

OpenStudy (anonymous):

Thank you very much!